OpenStudy (anonymous):

find d/dx of f^(-1) (x)

OpenStudy (anonymous):

what is the value of x?

OpenStudy (anonymous):

x/x + 1

OpenStudy (anonymous):

is f(x) = x/x + 1? or is f^-1(x)= x/x +1?

OpenStudy (anonymous):

d 2

OpenStudy (anonymous):

f^(-1) x = x/x+1

OpenStudy (anonymous):

well in any case, use the quotient rule where if you a function h defined as f/g then d/dx h = (f' * g - f * g') / g^2

OpenStudy (anonymous):

..............x f(x) = ------------- ..........x + 1 ........x y = -------- ......x + 1 ........y x = -------- ......y + 1 x(y + 1) = y xy + x = y xy - y = - x y(x - 1) = - x ...........x y = - --------- .........x - 1 .................x f⁻¹(x) = - ----------- answer// ................x - 1 bye..

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i got dat but it wasn't d ans

OpenStudy (anonymous):

you need to use the quotient rule next: x = f, x+1 = g d/dx x/x+1 = [1(x+1) - x(1)]/(x+1)^2 1/(x+1)^2

OpenStudy (anonymous):

yeah i got it now but it wasnt ur ans funinabox it was 1/(1-x)^2

OpenStudy (anonymous):

thank u all

OpenStudy (anonymous):

then you typed the problem wrong

OpenStudy (anonymous):

no it was correct

OpenStudy (anonymous):

then the answer is wrong 1/(1+x)^2 is the correct answer.

OpenStudy (anonymous):

but i got the marks

OpenStudy (anonymous):

i think i know what's up f(x) = x/x + 1 it's not f^-1(x) = x/x + 1 as was stated that way, when you do find the inverse, you get x/1-x, which would give you the derivative you posted.